I don't understand algebra at all and I understand slopes and lines even less! What is the equation for a line parallel to the line y=3x+4 and passes through the point (5, -3)

Thanks

Found 2 solutions by stanbon, josmiceli:
Answer by stanbon(75887)

What is the equation for a line parallel to the line y=3x+4 and passes through the point (5, -3)
-------------------
The given equation has a slope of 3.
Any parallel line must have slope of 3.
Form: y = mx+b
Substitute to solve for "b":
-3 = 3*5 + b
b = -3-15 = -18
EQUATION:
y = 3x-18
=============
Cheers,
Stan H.

Answer by josmiceli(19441)

You're wasting your time unless you get some really good
one-on-one help, because you have to go all the way back
to the most simple, basic stuff that you don't understand
and then build up from there. If you do that, and always
admit it when you don't know something, you will succeed.